# A Cross-Scene Benchmark for Open-World Drone Active Tracking

### v1

@@ C.2 More Details
Data Augmentation. In our two-stage curriculum learning process, we employ identical data augmentation. The flight altitude is selected from the interval $[13,22]\text{m}$, and the camera pitch angle is chosen from $[0.6,1.38]\text{rad}$. These parameters are consistent throughout each episode. Meanwhile, the drone’s initial orientation relative to the target fluctuates within the range $[-\pi,\pi]\text{rad}$, and the target’s initial position is set between $[-4.5,-2.5]\cup[2.5,4.5]\text{m}$.
Details on coordinate transformations. Given two planes $P_{0}:\hat{n_{0}}\textbf{x}^{T}+D_{0}=0$ and $P_{1}:\hat{n_{1}}\textbf{x}^{T}+D_{1}=0$, along with the HTM $T_{01}$ from $P_{0}$ to $P_{1}$. The transformation $T_{01}$ is defined as
@@
[…]
$\small T_{ct}=\begin{bmatrix}R_{ct}&t_{ct}\\ 0&1\\ \end{bmatrix},$ (30)
[…]
$\small T_{tc}=\begin{bmatrix}R^{T}_{ct}&-R^{T}_{ct}t_{ct}\\ 0&1\\ \end{bmatrix}.$ (31)
Using Eq. 29 and the transformation matrix $T_{tc}$, the plane $G_{c}$ can be formulated as $P_{G_{c}}=(r_{3}^{T},-h+r_{3}^{T}R^{T}_{ct}t_{ct})$.
@@
$\small r_{d}=\left\{\begin{array}[]{lr}1,&t\in I\\ 0,&\text{otherwise},\end{array}\right.$ (32)
where $I$ represents the image range of the drone. This reward can be used to construct the experimental metric, Tracking Success Rate (TSR).
Training algorithm. For the training method of R-VAT, we choose to use PPO algorithm. PPO algorithm regulates the speed of gradient updates by constraining the magnitude of policy changes $r_{t}$, expressed as follows:
@@
where $\hat{A_{t}}$ is the advantage function, $\epsilon$ is the clip parameter, and $\beta$ is the entropy coefficient. The expression of $\hat{A_{t}}$ is computed as:
@@
where $T,\lambda,\delta_{t+l}$ are the data collection step, generalized advantage estimator (GAE) [50] discount factor and temporal difference error respectively. The optimization objective expression of the critic network $V$ is defined as:
@@ Appendix D Baselines
Active Object Tracking (AOT) [33]. In this paper, the agent learns to follow a fixed target-tracking trajectory using the A3C algorithm. In addition, the Agent uses the following reward function:
@@
where $d$ represents the optimal distance between the tracker and the target, $c$ is the maximum allowable distance between the tracker and the target, and $A$ denotes the maximum reward. In the original paper, $c=200$ and $A=1.0$. During our replication, we set $A=1.0$, but due to the drone’s camera being tilted downward, a value of $c=200$ would far exceed the camera’s field of view, which is unrealistic. Therefore, we modify the parameter $c$ to be the maximum offset distance that keeps the target within the image frame, i.e., $c=9$.
Scene citystreet desert village downtown lake farmland
Steps (M) 19.2 13.4 21.3 19.8 9.9 9.2
Scene citystreet desert village downtown lake farmland
T (M) 10.0 6.2 8.0 10.3 5.6 4.1
Table 11: Transition steps across different scenes.
D-VAT[18]. In this approach, the agent uses an asymmetric Actor-Critic network structure and the soft actor-critic learning method [23] to accomplish the task of drone tracking another drone. In the actual comparative experiments, we convert it from a continuous action space to a discrete action space, referring to [10]. Additionally, the method uses the following reward function.

### v2

@@ C.2 More Details
Domain Randomization. While simpler settings facilitate the agent’s learning of task objectives, they also heighten the risk of the agent rapidly converging to a suboptimal action distribution, undermining the exploration process. Consequently, implementing domain randomization is essential. This is achieved through the randomization of the drone’s initial position and orientation relative to the target, necessitating a broader range of actions to maximize rewards. Moreover, to enhance the agent’s spatial perception ability, randomization is also introduced in its gimbal pitch angle.
In our two-stage curriculum learning process, we employ identical domain randomization. The flight altitude is selected from the interval $13,22]\text{m}$, and the camera pitch angle is chosen from $0.6,1.38]\text{rad}$. These parameters are consistent throughout each episode. Meanwhile, the drone’s initial orientation relative to the target fluctuates within the range $\pi,\pi]\text{rad}$, and the target’s initial position is set between $4.5,-2.5]\cup[2.5,4.5]\text{m}$.
Details on coordinate transformations. Given two planes $P_{0}:\hat{n_{0}}\textbf{x}^{T}\!\!+\!\!D_{0}\!=\!0$ and $P_{1}:\hat{n_{1}}\textbf{x}^{T}\!\!+\!\!D_{1}\!=\!0$, along with the HTM $T_{01}$ from $P_{0}$ to $P_{1}$. The $T_{01}$ is defined as:
@@
Considering the ground plane $G_{w}:z\!=\!h$ in the world coordinate system $\{\text{w}\}$, with representation in the camera coordinate system $\{\text{c}\}$ denoted as $G_{c}\!:\!A_{g}x\!+\!B_{g}y\!+\!C_{g}z\!+\!D_{g}\!=\!0$, the vectors of these two planes are $P_{G_{w}}\!=\!(0,0,1,-h)$ and $P_{G_{c}}\!=\!(A_{g},B_{g},C_{g},D_{g})$.
Furthermore, from Table 9, we can obtain the HTM $\texttt{T}_{cw}$ from $\{\text{c}\}$ to $\{\text{w}\}$ defined as follows:
$\small T_{cw}=\begin{bmatrix}R_{cw}&t_{cw}\\ 0&1\\ \end{bmatrix}$, (12)
where $R_{cw}$ is the rotation matrix from $\{\text{c}\}$ to $\{\text{w}\}$, which can be expressed in row vector form as: $R_{cw}=[r_{1},r_{2},r_{3}]^{T}$. Therefore, the homogeneous transformation matrix (HTM) $T_{wc}$, which represents the transformation from the world coordinate system $\{w\}$ to the camera coordinate system $\{c\}$, can be expressed as follows:
$\small T_{wc}=\begin{bmatrix}R^{T}_{cw}&-R^{T}_{cw}t_{cw}\\ 0&1\\ \end{bmatrix}$. (13)
Using Eq. 11 and the matrix $T_{wc}$, the plane $G_{c}$ can be formulated as $P_{G_{c}}=(r_{3}^{T},-h+r_{3}^{T}R^{T}_{cw}t_{cw})$.
Privileged knowledge available for Drone Agent. During training in the simulator, the drone agent has access to additional information (e.g., the precise location of the target). However, during testing and real-world deployment, such privileged knowledge is not available.
Table 12: The detailed results of comparison experiments on CRCR metric.
Within / Cross Scene Cross Domain
Train: citystreet
citystreet desert village downtown lake farmland night foggy snow
AOT $_{\pm 3} 49±94$$_{\pm 9}$ $45_{\pm 5}$ $49_{\pm 3}$ $48_{\pm 3}$ $48_{\pm 3}$ $49_{\pm 4}$ $49_{\pm 3}$ $49_{\pm 3}$
GC-VAT $\textbf{279}_{\pm 110}$ $\textbf{129}_{\pm 112}$ $\textbf{153}_{\pm 119}$ $\textbf{135}_{\pm 109}$ $\textbf{112}_{\pm 92}$ $\textbf{191}_{\pm 122}$ $\textbf{257}_{\pm 126}$ $\textbf{316}_{\pm 84}$ $\textbf{202}_{\pm 119}$
Train: desert
citystreet desert village downtown lake farmland night foggy snow
AOT $9_{\pm 0}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 0}$ $9_{\pm 0}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$
GC-VAT $\textbf{278}_{\pm 111}$ $\textbf{307}_{\pm 124}$ $\textbf{305}_{\pm 94}$ $\textbf{119}_{\pm 110}$ $\textbf{170}_{\pm 139}$ $\textbf{275}_{\pm 121}$ $\textbf{182}_{\pm 131}$ $\textbf{307}_{\pm 124}$ $\textbf{307}_{\pm 97}$
Train: village
citystreet desert village downtown lake farmland night foggy snow
D-VAT $46_{\pm 8}$ $5_{\pm 9} 44$$4_{\pm 8}$ $69_{\pm 42}$ $45_{\pm 8}$ $5_{\pm 3} 44$$_{\pm 8} 44$$4_{\pm 8}$ $43_{\pm 8}$
GC-VAT $\textbf{234}_{\pm 122}$ $\textbf{160}_{\pm 139}$ $\textbf{239}_{\pm 134}$ $\textbf{93}_{\pm 102}$ $\textbf{153}_{\pm 115}$ $\textbf{140}_{\pm 118}$ $\textbf{257}_{\pm 122}$ $\textbf{257}_{\pm 120}$ $\textbf{114}_{\pm 115}$
Train: downtown
citystreet desert village downtown lake farmland night foggy snow
D-VAT $8_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 2}$
GC-VAT $\textbf{209}_{\pm 131}$ $\textbf{184}_{\pm 136}$ $\textbf{202}_{\pm 129}$ $\textbf{203}_{\pm 119}$ $\textbf{189}_{\pm 93}$ $\textbf{223}_{\pm 114}$ $\textbf{167}_{\pm 135}$ $\textbf{165}_{\pm 126}$ $\textbf{178}_{\pm 125}$
Train: lake
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{112}_{\pm 86}$ $\textbf{144}_{\pm 110}$ $\textbf{203}_{\pm 133}$ $\textbf{143}_{\pm 134}$ $\textbf{181}_{\pm 116}$ $\textbf{214}_{\pm 111}$ $\textbf{190}_{\pm 129}$ $\textbf{168}_{\pm 110}$ $\textbf{99}_{\pm 67}$
Train: farmland
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{162}_{\pm 89}$ $\textbf{170}_{\pm 125}$ $\textbf{237}_{\pm 128}$ $\textbf{81}_{\pm 71}$ $\textbf{159}_{\pm 119}$ $\textbf{243}_{\pm 117}$ $\textbf{253}_{\pm 109}$ $\textbf{245}_{\pm 117}$ $\textbf{168}_{\pm 105}$
Table 13: The detailed results of comparison experiments on TSRTSR metric.
Within / Cross Scene Cross Domain
Train: citystreet
citystreet desert village downtown lake farmland night foggy snow
AOT $0.25_{\pm 0.02}$ $4_{\pm 0.03} 0.22±0$$2_{\pm 0.03}$ $0.25_{\pm 0.02}$ $0.23_{\pm 0.03}$ $0.24_{\pm 0.01}$ $0.25_{\pm 0.02}$ $0.25_{\pm 0.02}$ $0.24_{\pm 0.02}$
GC-VAT $\textbf{0.80}_{\pm 0.30}$ $\textbf{0.54}_{\pm 0.32}$ $\textbf{0.50}_{\pm 0.32}$ $\textbf{0.45}_{\pm 0.30}$ $\textbf{0.44}_{\pm 0.24}$ $\textbf{0.66}_{\pm 0.27}$ $\textbf{0.72}_{\pm 0.29}$ $\textbf{0.93}_{\pm 0.14}$ $\textbf{0.79}_{\pm 0.24}$
Train: desert
citystreet desert village downtown lake farmland night foggy snow
AOT $06_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.01} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.01} 0.0$$6_{\pm 0.01} 0.0$$06_{\pm 0.01}$
GC-VAT $\textbf{0.73}_{\pm 0.31}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{0.87}_{\pm 0.19}$ $\textbf{0.38}_{\pm 0.32}$ $\textbf{0.56}_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.25}$ $\textbf{0.57}_{\pm 0.31}$ $\textbf{0.86}_{\pm 0.28}$ $\textbf{0.86}_{\pm 0.22}$
Train: village
citystreet desert village downtown lake farmland night foggy snow
D-VAT $0.23_{\pm 0.04}$ $3_{\pm 0.04} 0.22±0$$2_{\pm 0.05}$ $0.31_{\pm 0.14}$ $4_{\pm 0.06} 0.22±0$$2_{\pm 0.01}$ $2_{\pm 0.04} 0.22±0$$2_{\pm 0.05}$ $0.23_{\pm 0.05}$
GC-VAT $\textbf{0.72}_{\pm 0.28}$ $\textbf{0.51}_{\pm 0.34}$ $\textbf{0.73}_{\pm 0.32}$ $\textbf{0.46}_{\pm 0.29}$ $\textbf{0.59}_{\pm 0.33}$ $\textbf{0.48}_{\pm 0.31}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.40}_{\pm 0.29}$
Train: downtown
citystreet desert village downtown lake farmland night foggy snow
D-VAT $05_{\pm 0.00} 0.0$$5_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.01} 0.0$$5_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.01} 0.0$$6_{\pm 0.00} 0.0$$06_{\pm 0.00}$
GC-VAT $\textbf{0.77}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.67}_{\pm 0.29}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.49}_{\pm 0.29}$ $\textbf{0.63}_{\pm 0.33}$ $\textbf{0.58}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.29}$ $\textbf{0.64}_{\pm 0.28}$
Train: lake
citystreet desert village downtown lake farmland night foggy snow
D-VAT $0.25_{\pm 0.04}$ $0.23_{\pm 0.04}$ $0.23_{\pm 0.05}$ $0.30_{\pm 0.15}$ $6_{\pm 0.06} 0.22±0$$2_{\pm 0.01}$ $0.26_{\pm 0.06}$ $0.26_{\pm 0.06}$ $0.25_{\pm 0.06}$
GC-VAT $\textbf{0.43}_{\pm 0.25}$ $\textbf{0.47}_{\pm 0.30}$ $\textbf{0.64}_{\pm 0.31}$ $\textbf{0.43}_{\pm 0.28}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{0.59}_{\pm 0.30}$ $\textbf{0.59}_{\pm 0.39}$ $\textbf{0.62}_{\pm 0.32}$ $\textbf{0.41}_{\pm 0.24}$
Train: farmland
citystreet desert village downtown lake farmland night foggy snow
AOT $0.24_{\pm 0.02}$ $4_{\pm 0.04} 0.22±0$$2_{\pm 0.03}$ $0.25_{\pm 0.02}$ $0.24_{\pm 0.02}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$
D-VAT $07_{\pm 0.01} 0.0$$7_{\pm 0.01} 0.0$$7_{\pm 0.00} 0.0$$8_{\pm 0.01} 0.0$$7_{\pm 0.00} 0.0$$7_{\pm 0.00} 0.0$$8_{\pm 0.00} 0.0$$7_{\pm 0.00} 0.0$$08_{\pm 0.00}$
GC-VAT $\textbf{0.48}_{\pm 0.24}$ $\textbf{0.59}_{\pm 0.34}$ $\textbf{0.72}_{\pm 0.26}$ $\textbf{0.33}_{\pm 0.20}$ $\textbf{0.58}_{\pm 0.28}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{0.67}_{\pm 0.32}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.51}_{\pm 0.28}$
@@
$\small r_{d}=\left\{\begin{array}[]{lr}1,&t\in\mathcal{I}\\ 0,&\text{otherwise}\end{array}\right$., (14)
where $\mathcal{I}$ represents the image range. This reward can be used to construct the metric, Tracking Success Rate (TSR).
Figure 13: Schematic diagram of reward curves on DAT scenes.
Training algorithm. For the training method of GC-VAT, we choose to use PPO algorithm. PPO algorithm regulates the speed of gradient updates by constraining the magnitude of policy changes $r_{t}$, expressed as follows:
@@
where $\hat{A_{t}}$ is the advantage function, $\epsilon$ is the clip parameter, and $\beta$ is the entropy coefficient. The expression of $\hat{A_{t}}$ is:
@@
where $T,\lambda,\delta_{t+l}$ are the data collection step, generalized advantage estimator (GAE) gae discount factor and temporal difference error respectively. The optimization objective expression of the critic network $V$ is defined as:
@@
Table 14: Effectiveness of CBT strategy on the DAT benchmark, results from CRCR metric.
Within / Cross Scene Cross Domain
Train: citystreet
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{279}_{\pm 110}$ $\textbf{129}_{\pm 112}$ $\textbf{153}_{\pm 119}$ $\textbf{135}_{\pm 109}$ $\textbf{112}_{\pm 92}$ $\textbf{191}_{\pm 122}$ $\textbf{257}_{\pm 126}$ $\textbf{316}_{\pm 84}$ $\textbf{202}_{\pm 119}$
Train: desert
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{278}_{\pm 111}$ $\textbf{307}_{\pm 124}$ $\textbf{305}_{\pm 94}$ $119_{\pm 110}$ $170_{\pm 139}$ $\textbf{275}_{\pm 121}$ $182_{\pm 131}$ $\textbf{307}_{\pm 124}$ $\textbf{307}_{\pm 97}$
Train: village
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{234}_{\pm 122}$ $160_{\pm 139}$ $239_{\pm 134}$ $\textbf{93}_{\pm 102}$ $\textbf{153}_{\pm 115}$ $140_{\pm 118}$ $257_{\pm 122}$ $\textbf{257}_{\pm 120}$ $\textbf{114}_{\pm 115}$
Train: downtown
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{209}_{\pm 131}$ $\textbf{184}_{\pm 136}$ $\textbf{202}_{\pm 129}$ $\textbf{203}_{\pm 119}$ $\textbf{189}_{\pm 93}$ $\textbf{223}_{\pm 114}$ $\textbf{167}_{\pm 135}$ $\textbf{165}_{\pm 126}$ $\textbf{178}_{\pm 125}$
Train: lake
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $112_{\pm 86}$ $\textbf{144}_{\pm 110}$ $\textbf{203}_{\pm 133}$ $\textbf{143}_{\pm 134}$ $181_{\pm 116}$ $\textbf{214}_{\pm 111}$ $\textbf{190}_{\pm 129}$ $168_{\pm 110}$ $99_{\pm 67}$
Train: farmland
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{162}_{\pm 89}$ $\textbf{170}_{\pm 125}$ $\textbf{237}_{\pm 128}$ $\textbf{81}_{\pm 71}$ $\textbf{159}_{\pm 119}$ $\textbf{243}_{\pm 117}$ $\textbf{253}_{\pm 109}$ $\textbf{245}_{\pm 117}$ $\textbf{168}_{\pm 105}$
Table 15: Effectiveness of CBT strategy on the DAT benchmark, results from TSRTSR metric.
Within / Cross Scene Cross Domain
Train: citystreet
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{0.80}_{\pm 0.30}$ $\textbf{0.54}_{\pm 0.32}$ $\textbf{0.50}_{\pm 0.32}$ $\textbf{0.45}_{\pm 0.30}$ $\textbf{0.44}_{\pm 0.24}$ $\textbf{0.66}_{\pm 0.27}$ $\textbf{0.72}_{\pm 0.29}$ $\textbf{0.93}_{\pm 0.14}$ $\textbf{0.79}_{\pm 0.24}$
Train: desert
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $0.73_{\pm 0.31}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{0.87}_{\pm 0.19}$ $0.38_{\pm 0.32}$ $0.56_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.25}$ $0.57_{\pm 0.31}$ $\textbf{0.86}_{\pm 0.28}$ $\textbf{0.86}_{\pm 0.22}$
Train: village
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $0.72_{\pm 0.28}$ $0.51_{\pm 0.34}$ $0.73_{\pm 0.32}$ $\textbf{0.46}_{\pm 0.29}$ $\textbf{0.59}_{\pm 0.33}$ $0.48_{\pm 0.31}$ $0.71_{\pm 0.32}$ $\textbf{0.71}_{\pm 0.32}$ $0.40_{\pm 0.29}$
Train: downtown
citystreet desert village downtown lake farmland night foggy snow
w/o CBT $0.29_{\pm 0.04}$ $0.27_{\pm 0.03}$ $0.27_{\pm 0.03}$ $0.33_{\pm 0.06}$ $0.28_{\pm 0.03}$ $0.27_{\pm 0.01}$ $3_{\pm 0.06} 0.33±0$$3_{\pm 0.06}$ $0.33_{\pm 0.06}$
GC-VAT $\textbf{0.77}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.67}_{\pm 0.29}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.49}_{\pm 0.29}$ $\textbf{0.63}_{\pm 0.33}$ $\textbf{0.58}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.29}$ $\textbf{0.64}_{\pm 0.28}$
Train: lake
citystreet desert village downtown lake farmland night foggy snow
[…]
GC-VAT $0.43_{\pm 0.25}$ $0.47_{\pm 0.30}$ $\textbf{0.64}_{\pm 0.31}$ $0.43_{\pm 0.28}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{0.59}_{\pm 0.30}$ $0.59_{\pm 0.39}$ $0.62_{\pm 0.32}$ $\textbf{0.41}_{\pm 0.24}$
Train: farmland
citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{0.48}_{\pm 0.24}$ $\textbf{0.59}_{\pm 0.34}$ $\textbf{0.72}_{\pm 0.26}$ $\textbf{0.33}_{\pm 0.20}$ $\textbf{0.58}_{\pm 0.28}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{0.67}_{\pm 0.32}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.51}_{\pm 0.28}$
@@ Appendix D Baselines
Table 16: Effectiveness of reward design on the DAT benchmark, results from CRCR metric.
Within / Cross Scene Cross Domain
Train: citystreet citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{279}_{\pm 110}$ $\textbf{129}_{\pm 112}$ $\textbf{153}_{\pm 119}$ $\textbf{135}_{\pm 109}$ $\textbf{112}_{\pm 92}$ $\textbf{191}_{\pm 122}$ $\textbf{257}_{\pm 126}$ $\textbf{316}_{\pm 84}$ $\textbf{202}_{\pm 119}$
Train: desert citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{278}_{\pm 111}$ $\textbf{307}_{\pm 124}$ $\textbf{305}_{\pm 94}$ $\textbf{119}_{\pm 110}$ $\textbf{170}_{\pm 139}$ $\textbf{275}_{\pm 121}$ $\textbf{182}_{\pm 131}$ $\textbf{307}_{\pm 124}$ $\textbf{307}_{\pm 97}$
Train: village citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{234}_{\pm 122}$ $\textbf{160}_{\pm 139}$ $\textbf{239}_{\pm 134}$ $\textbf{93}_{\pm 102}$ $\textbf{153}_{\pm 115}$ $\textbf{140}_{\pm 118}$ $\textbf{257}_{\pm 122}$ $\textbf{257}_{\pm 120}$ $\textbf{114}_{\pm 115}$
Train: downtown citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{209}_{\pm 131}$ $\textbf{184}_{\pm 136}$ $\textbf{202}_{\pm 129}$ $\textbf{203}_{\pm 119}$ $\textbf{189}_{\pm 93}$ $\textbf{223}_{\pm 114}$ $\textbf{167}_{\pm 135}$ $\textbf{165}_{\pm 126}$ $\textbf{178}_{\pm 125}$
Train: lake citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $_{\pm 3} 11±11$$_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 2}$ $9_{\pm 0}$ $8_{\pm 0}$ $9_{\pm 0}$ $10_{\pm 1}$ $8_{\pm 1}$
GC-VAT $\textbf{112}_{\pm 86}$ $\textbf{144}_{\pm 110}$ $\textbf{203}_{\pm 133}$ $\textbf{143}_{\pm 134}$ $\textbf{181}_{\pm 116}$ $\textbf{214}_{\pm 111}$ $\textbf{190}_{\pm 129}$ $\textbf{168}_{\pm 110}$ $\textbf{99}_{\pm 67}$
Train: farmland citystreet desert village downtown lake farmland night foggy snow
GC-VAT $\textbf{162}_{\pm 89}$ $\textbf{170}_{\pm 125}$ $\textbf{237}_{\pm 128}$ $\textbf{81}_{\pm 71}$ $\textbf{159}_{\pm 119}$ $\textbf{243}_{\pm 117}$ $\textbf{253}_{\pm 109}$ $\textbf{245}_{\pm 117}$ $\textbf{168}_{\pm 105}$
Table 17: Effectiveness of reward design on the DAT benchmark, results from TSRTSR metric.
Within / Cross Scene Cross Domain
Train: citystreet
citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $06_{\pm 0.00} 0.0$$5_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.01} 0.0$$06_{\pm 0.00}$
GC-VAT $\textbf{0.80}_{\pm 0.30}$ $\textbf{0.54}_{\pm 0.32}$ $\textbf{0.50}_{\pm 0.32}$ $\textbf{0.45}_{\pm 0.30}$ $\textbf{0.44}_{\pm 0.24}$ $\textbf{0.66}_{\pm 0.27}$ $\textbf{0.72}_{\pm 0.29}$ $\textbf{0.93}_{\pm 0.14}$ $\textbf{0.79}_{\pm 0.24}$
Train: desert
citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $06_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.01} 0.0$$06_{\pm 0.00}$ $10_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$9_{\pm 0.01} 0.0$$06_{\pm 0.00}$
GC-VAT $\textbf{0.73}_{\pm 0.31}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{0.87}_{\pm 0.19}$ $\textbf{0.38}_{\pm 0.32}$ $0.56_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.25}$ $\textbf{0.57}_{\pm 0.31}$ $\textbf{0.86}_{\pm 0.28}$ $\textbf{0.86}_{\pm 0.22}$
Train: village
citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $06_{\pm 0.01} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.01} 0.0$$5_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$5_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$06_{\pm 0.00}$
GC-VAT $\textbf{0.72}_{\pm 0.28}$ $\textbf{0.51}_{\pm 0.34}$ $\textbf{0.73}_{\pm 0.32}$ $\textbf{0.46}_{\pm 0.29}$ $\textbf{0.59}_{\pm 0.33}$ $\textbf{0.48}_{\pm 0.31}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.40}_{\pm 0.29}$
Train: downtown
citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $06_{\pm 0.01} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$5_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$06_{\pm 0.00}$
GC-VAT $\textbf{0.77}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.67}_{\pm 0.29}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.49}_{\pm 0.29}$ $\textbf{0.63}_{\pm 0.33}$ $\textbf{0.58}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.29}$ $\textbf{0.64}_{\pm 0.28}$
Train: lake
citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $10_{\pm 0.01} 0.0$$9_{\pm 0.01} 0.0$$7_{\pm 0.00} 0.0$$6_{\pm 0.01} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$8_{\pm 0.00} 0.0$$06_{\pm 0.00}$
GC-VAT $\textbf{0.43}_{\pm 0.25}$ $\textbf{0.47}_{\pm 0.30}$ $\textbf{0.64}_{\pm 0.31}$ $\textbf{0.43}_{\pm 0.28}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{0.59}_{\pm 0.30}$ $\textbf{0.59}_{\pm 0.39}$ $\textbf{0.62}_{\pm 0.32}$ $\textbf{0.41}_{\pm 0.24}$
Train: farmland
citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $06_{\pm 0.00} 0.0$$5_{\pm 0.00} 0.0$$5_{\pm 0.00} 0.0$$6_{\pm 0.01} 0.0$$5_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$6_{\pm 0.00} 0.0$$06_{\pm 0.00}$
GC-VAT $\textbf{0.48}_{\pm 0.24}$ $\textbf{0.59}_{\pm 0.34}$ $\textbf{0.72}_{\pm 0.26}$ $\textbf{0.33}_{\pm 0.20}$ $\textbf{0.58}_{\pm 0.28}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{0.67}_{\pm 0.32}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.51}_{\pm 0.28}$
Figure 14: Qualitative results on images from the car6 video sequence. Arrows link data points to the visualization of associated scenarios.
Active Object Tracking (AOT) e2e . In this paper, the agent learns to follow a fixed target-tracking trajectory using A3C. In addition, the agent uses the following reward:
@@
where $d$ represents the optimal distance between the tracker and the target, $c$ is the maximum allowable distance, and $A$ denotes the maximum reward. In the original paper, $c=200$ and $A=1.0$. During our replication, we set $A=1.0$, but due to the drone’s camera being tilted downward, a value of $c=200$ would far exceed the camera’s field of view, which is unrealistic. Therefore, we modify the parameter $c$ to be the maximum offset distance that keeps the target within the image, i.e., $c=9$.
D-VATdvat . In this approach, the agent uses an asymmetric Actor-Critic network structure and the soft actor-critic learning method sac to accomplish the task of drone tracking another drone. In the actual comparative experiments, we convert it from a continuous action space to a discrete action space, referring to sacd . Additionally, the method uses the following reward function.